# Designing Neural Synthesizers for Low Latency Interaction

### v1

@@ II-B Interactive Neural Audio Synthesis
NAS is a technique that employs Deep Neural Networks (DNNs) as learned audio synthesizers. Neural synthesis has its origins in offline generation; however, increased computation capabilities have enabled real-time inference, which has opened the doors for NAS use in live performance systems with control schemes that span from MIDI inputs [2, 15] to MIR features [16], latent spaces [17, 18], and other audio signals [4, 19]. Despite this, latency is typically only addressed after models are trained [8]. A common solution in neural audio plug-in development is to report latency to the audio workstation for it to be compensated by re-aligning other audio tracks, which is incompatible with our live use case.
@@ III-C Case study: RAVE
Waveform Generation RAVE generates audio samples with a waveform synthesizer and a noise generator. The waveform synthesizer projects the decoder’s output and generates a loudness envelope and multiband signal (with tanh activation) that is inverted back into 2048 samples using PQMF. The noise generator is based on a strided convolutional stack that re-compresses the decoder’s output using a ratio of 1024 and then performs FFT convolution on white noise to generate another multiband signal that is summed with the waveform synthesizer before PQMF inversion.
@@ IV-B Latency and RTF Evaluation
We design a basic C++ application that loads the scripted models and times their execution time for different audio block sizes. The models are executed 1100 times, with the first 100 executions employed as a ”warm-up” for the OS scheduler to find an optimal way to handle the inference workload, and the rest registered for computing statistics. Then, our application calculates the RTF of each run taking 44.1 kHz as a reference sampling rate and reports the mean and standard deviation. We perform all measurements on a MacBook with an M1 Pro CPU and 16 GB RAM, running MacOS Sonoma 14.3.
@@ IV-F Low Latency Implementation
We implement three different causal models in our architecture: a version of RAVE v1 without noise generator (c2048_r10), a low latency and high-capacity model (c128_r05_p40), and BRAVE. We achieve the fastest implementation in our reference M1 Pro CPU with the STL backend of RTNeural, using Apple Clang 15. Table V compares the RTF between the RTNeural and Libtorch implementations, showing that the latter can execute BRAVE comfortably at the lowest block size and latency, unlike Libtorch. However, it does not improve its RTF with longer block sizes for any of the models, with Libtorch becoming a more efficient option for longer block sizes. We suspect Libtorch may employ computing algorithms optimized for different block sizes. We deploy our models in a proof-of-concept audio plugin (see the accompanying webpage).
@@ V-F Discussion of Results
We believe this approach can work for other low-latency interactive NAS tasks, provided all modules within the network are designed to support such latent trajectories. For instance, a low-latency re-design of RAVE’s noise generator for low compression ratio operation could improve BRAVE’s audio quality. One possibility would be the addition of low-latency, differentiable IIR filters [78] controlled by an additional decoder submodule with an appropriate receptive field. We expect our design experience can guide other researchers in the design of complex low-latency NAS algorithms, while leveraging recent inference tools that can simplify implementation [66, 65].

### v2

# Designing Neural Synthesizers for Low-Latency Interaction

@@ II-B Interactive Neural Audio Synthesis
NAS is a technique that employs Deep Neural Networks (DNNs) as learned audio synthesizers. Neural synthesis has its origins in offline generation; however, increased computation capabilities have enabled real-time inference, which has opened the doors for NAS use in live performance systems with control schemes that span from MIDI inputs [2, 15] to Music Information Retrieval features [16], latent spaces [17, 18], and other audio signals [4, 19]. Despite this, latency is typically only addressed after models are trained [8]. A common solution in neural audio plug-in development is to report latency to the audio workstation for it to be compensated by re-aligning other audio tracks, which is incompatible with our live use case.
@@ III-C Case study: RAVE
Waveform Generation RAVE generates audio samples with a waveform synthesizer and a noise generator. The waveform synthesizer projects the decoder’s output and generates a loudness envelope and multiband signal (with tanh activation) that is inverted back into 2048 samples using PQMF. The noise generator is based on a strided convolutional stack that re-compresses the decoder’s output using a ratio of 1024 and then performs Fast Fourier Transform (FFT) convolution on white noise to generate another multiband signal that is summed with the waveform synthesizer before PQMF inversion.
@@ IV-B Latency and RTF Evaluation
We design a basic C++ application that loads the scripted models and times their execution time for different audio block sizes. The models are executed 1100 times, with the first 100 executions employed as a ”warm-up” for the operating system scheduler to find an optimal way to handle the inference workload, and the rest registered for computing statistics. Then, our application calculates the RTF of each run taking 44.1 kHz as a reference sampling rate and reports the mean and standard deviation. We perform all measurements on a MacBook with an M1 Pro CPU and 16 GB RAM, running MacOS Sonoma 14.3.
@@ IV-F Low Latency Implementation
We implement three different causal models in our architecture: a version of RAVE v1 without noise generator (c2048_r10), a low latency and high-capacity model (c128_r05_p40), and BRAVE. We achieve the fastest implementation in our reference M1 Pro CPU with the Standard Template Library backend of RTNeural, using Apple Clang 15. Table V compares the RTF between the RTNeural and Libtorch implementations, showing that the latter can execute BRAVE comfortably at the lowest block size and latency, unlike Libtorch. However, it does not improve its RTF with longer block sizes for any of the models, with Libtorch becoming a more efficient option for longer block sizes. We suspect Libtorch may employ computing algorithms optimized for different block sizes. We deploy our models in a proof-of-concept audio plugin (see the accompanying webpage).
@@ V-F Discussion of Results
We believe this approach can work for other low-latency interactive NAS tasks, provided all modules within the network are designed to support such latent trajectories. For instance, a low-latency re-design of RAVE’s noise generator for low compression ratio operation could improve BRAVE’s audio quality. One possibility would be the addition of low-latency, differentiable Infinite Impulse Response filters [77] controlled by an additional decoder submodule with an appropriate receptive field. We expect our design experience can guide other researchers in the design of complex low-latency NAS algorithms, while leveraging recent inference tools that can simplify implementation [65, 64].